A uniform electric field E = 5000 N/C passing through a flat square area A = 2m2. Determine the electric flux. Note: The angle between the electric field direction and a line drawn perpendicular to the area is 60 degree.

1
Expert's answer
2022-04-11T12:08:50-0400

The electric flux is given by

"\phi=EA\cos\theta""\phi=5000*2*\cos60^\circ=5000\:\rm N\cdot m^2/C"